OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by three times cosine of pi divided by five plus sine of pi divided by three times sine of pi divided by five.

OpenStudy (anonymous):

@mathmath333 please please please help?

OpenStudy (mathmath333):

can u write the equation

OpenStudy (anonymous):

\[\cos (\pi/3) * \cos (\pi/5) + \sin (\pi/3) * \sin(\pi/5)\]

OpenStudy (mathmath333):

u can use the formula here \(\large\tt \color{black}{cos(A-B)=cosA\times cosB+sinA\times sinB}\) \(\large\tt \color{black}{cos(\frac{\pi}{3})\times cos(\frac{\pi}{5})+sin(\frac{\pi}{3})\times sin(\frac{\pi}{5})}\) \(\large\tt \color{black}{=cos(\frac{\pi}{3}-\frac{\pi}{5})}\) \(\large\tt \color{black}{=cos(\frac{5\pi-3\pi}{15})}\) \(\large\tt \color{black}{=cos(\frac{2\pi}{15})}\)

OpenStudy (anonymous):

Thank you. I had that answer. And the math. I just thought I was wrong. I appreciate it.